OpenStudy (anonymous):

simplify sqrt(45/20)

OpenStudy (amistre64):

9.5 --- .... 3/2 4.5

OpenStudy (dumbcow):

break it into 2 parts sqrt(45)/sqrt(20) then try to factor out perfect squares from each one 45 = 9*5 20 = 4*5 sqrt(9)*sqrt(5) / sqrt(4)*sqrt(5) = 3sqrt(5)/2sqrt(5) =3/2